OpenStudy (anonymous):

-4 square root x + 9 =20 How do i solve this step by step? so confused

OpenStudy (anonymous):

\[-4 \sqrt{x+9} = 20\]Is this what you mean?

OpenStudy (anonymous):

The first step would be to undo the multiplication by using division. \[\sqrt{x+9}=\frac{ 20 }{ -4 }\]\[\sqrt{x+9} = -5\]

OpenStudy (anonymous):

nw square both sides

OpenStudy (anonymous):

x + 9 = 25 x = 16

OpenStudy (anonymous):

Then to get rid of the square root, you square both sides:\[(\sqrt{x+9})^{2} = (-5)^{2}\]\[x+9 = 25\]\[x=16\]

OpenStudy (anonymous):

Remember, if you put "-5^2 into your calculator, you will get -25, because the calculator does the 5^2 part first, then multiplies the result by -1. So you need to put in (-5)^2 (with the brackets)!

OpenStudy (theviper):

\[-4 \sqrt{x+9} = 20\]\[\sqrt{x+9}=\frac{20}{-4}=-5\]\[x+9=25\implies x=25-9=16\]